Working with your cousin, you can refinish a table in 6 hours. Working
alone, your cousin can complete the job in 9 hours. How long would
it take you to refinish the table?
alone?

Answer by richwmiller(17219)

6/x+6/y=1
y=9
6/x+6/9=1
6/x+2/3=1
6/x=1/3
18=x
18 hours